# Joint Activity Detection and Channel Estimation for Massive Connectivity: Where Message Passing Meets Score-Based Generative Priors

### v1

@@ III-A Module A: Linear Estimator
Following Turbo-CS [25], we start by initializing the message from variable node $\mathbf{x}_{m}$ to factor node $p(\mathbf{y}_{m}|\mathbf{x}_{m})$ as $\mathcal{M}_{\mathbf{x}_{m}\rightarrow f_{\mathbf{y}_{m}}}(\mathbf{x}_{m})=\mathcal{CN}(\mathbf{x}_{m};\mathbf{x}_{m,\mathsf{A}}^{\mathtt{pri}},v_{m,\mathsf{A}}^{\mathtt{pri}}\mathbf{I}_{NK})$. The belief of $\mathbf{x}_{m}$ on factor node $p(\mathbf{y}_{m}|\mathbf{x}_{m})$ is expressed as
@@
The message from factor node $p(\mathbf{y}_{m}|\mathbf{x}_{m})$ to variable node $\mathbf{x}_{m}$, also known as the extrinsic message, is calculated based on the EP algorithm [48] as
@@ III-B Module B: Channel and Device Activity Denoisers
As shown in Fig. 2, the denoiser of $\mathbf{X}_{k}=\alpha_{k}\mathbf{H}_{k}$ can be further divided into two sub-modules: the denoiser of $\mathbf{H}_{k}$ and the denoiser of $\alpha_{k}$. The outputs from these two sub-modules are then combined at the factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ to produce the denoised $\mathbf{X}_{k}$. In this subsection, we formally establish the equivalence between the belief calculation at the factor node $p(\mathbf{H}_{k})$ and the MMSE denoising of $\mathbf{H}_{k}$ under AWGN observations. This equivalence plays a crucial role in the integration of score-based generative priors for efficient belief calculation.
Firstly, we derive the message from variable node $\mathbf{X}_{k}$ to factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$, denoted as $\mathcal{M}_{\mathbf{X}_{k}\rightarrow f_{\mathbf{X}_{k}}}(\mathbf{X}_{k})$, by examining the relationships in vector-matrix reorganization. We marginalize the message $\mathcal{M}_{f_{\mathbf{y}_{m}}\rightarrow\mathbf{x}_{m}}(\mathbf{x}_{m})=\mathcal{CN}\big{(}\mathbf{x}_{m};\mathbf{x}_{m,\mathsf{B}}^{\mathtt{pri}},v_{m,\mathsf{B}}^{\mathtt{pri}}\mathbf{I}_{NK}\big{)}$ from module A to obtain
@@
where $\mathbf{x}_{km}\triangleq\left[x_{k1m},\dots,x_{kNm}\right]^{\sf T}\in\mathbb{C}^{N}$ and $\mathbf{x}_{km,\mathsf{B}}^{\mathtt{pri}}\triangleq\big{[}x_{k1m,\mathsf{B}}^{\mathtt{pri}},\dots,x_{kNm,\mathsf{B}}^{\mathtt{pri}}\big{]}^{\sf T}\in\mathbb{C}^{N}$ represent the $k$-th blocks in $\mathbf{x}_{m}$ and $\mathbf{x}_{m,\mathsf{B}}^{\mathtt{pri}}$, respectively, i.e., $\mathbf{x}_{m}=\left[\mathbf{x}_{1m}^{\sf H},\dots,\mathbf{x}_{Km}^{\sf H}\right]^{\sf H}$ and $\mathbf{x}_{m,\mathsf{B}}^{\mathtt{pri}}=\big{[}(\mathbf{x}_{1m,\mathsf{B}}^{\mathtt{pri}})^{\sf H},\dots,(\mathbf{x}_{Km,\mathsf{B}}^{\mathtt{pri}})^{\sf H}\big{]}^{\sf H}$. After vector-matrix reorganization, the message from variable node $\mathbf{X}_{k}$ to factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ manifests as
@@ III-B1 Denoiser of $\mathbf{H}_{k}$
The message from factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ to variable node $\mathbf{H}_{k}$ is computed following the sum-product rule [47]:
@@
where $\mathbf{h}_{km}\triangleq\left[h_{k1m},\dots,h_{kNm}\right]^{\sf T}\in\mathbb{C}^{N}$ represents the $m$-th column of $\mathbf{H}_{k}$, i.e., $\mathbf{H}_{k}=\left[\mathbf{h}_{k1},\dots,\mathbf{h}_{kM}\right]$. The first term in (15) is a constant irrelevant to $\mathbf{H}_{k}$, making the normalization of $\mathcal{M}_{f_{\mathbf{X}_{k}}\rightarrow\mathbf{H}_{k}}(\mathbf{H}_{k})$ to be a pdf problematic. Note that the first and the second terms correspond to the $k$-th device being inactive and active, respectively. For the inactive case, it is unnecessary to pass the message to factor node $p(\mathbf{H}_{k})$ for denoising, as our focus is solely on estimating active channels. Therefore, we approximate $\mathcal{M}_{f_{\mathbf{X}_{k}}\rightarrow\mathbf{H}_{k}}(\mathbf{H}_{k})$ by only taking the second term into account, resulting in a Gaussian message of the form
@@
Based on the update rules of EP [48], the belief of $\mathbf{H}_{k}$ on factor node $p(\mathbf{H}_{k})$ is expressed as
@@
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\prod_{m}\mathcal{CN}\big{(}\mathbf{h}_{km};\mathbf{h}_{km}^{\mathtt{pri}},\tau_{m}^{\mathtt{pri}}\mathbf{I}_{N}\big{)},$ (19)
@@
The state evolution (SE) analysis [50] has shown that the error vector $\mathbf{h}_{km}^{\mathtt{pri}}-\mathbf{h}_{km}$ is i.i.d. Gaussian in the large system limit. In other words, module A decouples the linear mixing model (3) into a series of independent AWGN observations of the ground truth, i.e.,
@@
where $\mathbf{H}_{k}^{\mathtt{post}}\triangleq\left[\mathbf{h}_{k1}^{\mathtt{post}},\dots,\mathbf{h}_{kM}^{\mathtt{post}}\right]$. The posterior mean and variance in (22) and (23) are, however, by no means easy to calculate due to the complicated and high-dimensional integrals. We will detail the method for obtaining the posterior message in Sections IV and V.
The extrinsic message from variable node $\mathbf{H}_{k}$ to factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ is calculated as
@@ III-B2 Denoiser of $\alpha_{k}$
The belief of $\alpha_{k}$ on factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ is expressed as
@@
is the posterior estimate of the active probability of the $k$-th device. We set a threshold $\lambda^{\mathtt{thr}}$ to facilitate the binary decision of device activity, i.e.,
@@ III-B3 Denoiser of $\mathbf{X}_{k}$
Based on the update rules of EP [48], the belief of $\mathbf{X}_{k}$ on factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ is expressed as
@@
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\mathcal{M}_{\mathbf{H}_{k}\rightarrow f_{\mathbf{X}_{k}}}(\mathbf{H}_{k})p(\alpha_{k})\bigg{]}$
@@
Finally, we calculate the message from factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ to variable node $\mathbf{X}_{k}$ as
@@ IV-A Bridging Score Function and MMSE Denoising
where $\nabla_{\mathbf{H}_{k}^{\mathtt{pri}}}\log p(\mathbf{H}_{k}^{\mathtt{pri}})\in\mathbb{C}^{N\times M}$ is the gradient of the log-density function, commonly known as the (first-order) score function. Tweedie’s formula bridges MMSE denoising and the score of the perturbed data distribution through a simple equation, eliminating the need for calculating high-dimensional integrals. Specifically, the posterior mean can be viewed as the sum of the maximum likelihood estimator and a Bayes correction term. The formula can also be framed as “denoised result = noisy data - predicted noise” from a denoising perspective, in which the predicted noise corresponds to $-\tau^{\mathtt{pri}}\nabla_{\mathbf{H}_{k}^{\mathtt{pri}}}\log p(\mathbf{H}_{k}^{\mathtt{pri}})$.
@@
where $\nabla_{\mathbf{H}_{k}^{\mathtt{pri}}}^{2}\log p(\mathbf{H}_{k}^{\mathtt{pri}})\in\mathbb{C}^{NM\times NM}$ is the Hessian of the log-density function, referred to as the second-order score function. The first- and second-order Tweedie’s formulas are well-known results that can be proven using simple integration tricks [51, Appendix B] or by applying the properties of exponential families [24, Appendix A].
@@ IV-B First-Order Score Matching
Eqn. (46) cannot be directly used as the loss function since $\nabla_{\tilde{\mathbf{H}}}\log p(\tilde{\mathbf{H}})$ does not have an analytical form. The above formulation is shown equivalent to the denoising score matching objective [29, 15], given by
@@
The equivalence is in the sense that (46) and (47) share the same optimal solution $\mathbf{S}_{1}(\tilde{\mathbf{H}},\tau;\boldsymbol{\theta}^{\star})=\nabla_{\tilde{\mathbf{H}}}\log p(\tilde{\mathbf{H}})$, supposing that the network has sufficient representation capability. Substituting $\nabla_{\tilde{\mathbf{H}}}\log p(\tilde{\mathbf{H}}|\mathbf{H})=-\frac{\tilde{\mathbf{H}}-\mathbf{H}}{\tau}$ into (47) yields
@@
A common choice for the weighting factor is $\lambda_{1}(\tau_{i})=\tau_{i}^{2}$ [15], which empirically balances the loss term $\ell_{1}(\tau_{i};\boldsymbol{\theta})$ for different values of $\tau_{i}$.
@@ IV-C Second-Order Score Matching
The second-order score matching objective can also be expressed in an equivalent denoising score matching formulation [30] as
@@
where $\mathbf{b}(\mathbf{H},\tilde{\mathbf{H}},\tau)\triangleq\mathrm{vec}\left(\nabla_{\tilde{\mathbf{H}}}\log p(\tilde{\mathbf{H}})+\frac{\tilde{\mathbf{H}}-\mathbf{H}}{\tau}\right)$. Please refer to [24, Appendix B] for the detailed derivation. The objective in (51) relies on the ground truth of the first-order score function, which is in general unavailable. To address this, we replace $\nabla_{\tilde{\mathbf{H}}}\log p(\tilde{\mathbf{H}})$ with the learned model $\mathbf{S}_{1}(\tilde{\mathbf{H}},\tau;\boldsymbol{\theta})$ for efficient computation. Consequently, the objective in (51) modifies to
@@
where $\lambda_{2}(\tau_{i})$ is the weighting factor depending on $\tau_{i}$. Following [30], we choose $\lambda_{2}(\tau_{i})=\tau_{i}^{4}$.
@@
Figure 3: Diagram of the STMP-JADCE algorithm. The real and imaginary parts of the channel matrix are separated into two RGB channels when input into the score networks. In this figure, the first and the second devices are active, whereas the $K$-th device is inactive. Therefore, the input to the score networks takes the form of $\mathbf{H}_{k}^{\mathtt{pri}}=\mathbf{H}_{k}+\mathbf{W}_{k}$ for $k=1,2$, while $\mathbf{H}_{K}^{\mathtt{pri}}$ is a random noise. Note that, for brevity, we omitted the skip connections (dependences) from the prior messages to both the the posterior and extrinsic messages. Please refer to eqns. (55), (56), (26), and (27) for the rigorous relationships.
@@ V-A Overall Algorithm
where $s_{2,nm}(\mathbf{H}_{k}^{\mathtt{pri}},\tau^{\mathtt{pri}};\boldsymbol{\phi})$ denotes the $((n-1)M+m)$-th diagonal elements of $\mathbf{S}_{2}\big{(}\mathbf{H}_{k}^{\mathtt{pri}},\tau^{\mathtt{pri}};\boldsymbol{\phi}\big{)}$.
Fig. 3 demonstrates the integration of the two score networks into the TMP framework, resulting in the STMP-JADCE algorithm. STMP-JADCE features a hybrid structure in which the channel denoiser is realized by score networks, while the device activity denoiser and the linear estimator are derived analytically. In this way, our method can be easily applied across different device activity levels, pilot patterns, and SNRs without the need to retain the score networks. Moreover, the potential mismatch between the chosen prior of the device activity and the ground truth can be rectified by incorporating the expectation-maximization (EM) update [52]. We discuss the engineering tricks used in STMP-JADCE in the following.

### v2

@@ III-A Module A: Linear Estimator
Following Turbo-CS [26], we start by initializing the message from the variable node $\mathbf{x}_{m}$ to the factor node $p(\mathbf{y}_{m}|\mathbf{x}_{m})$ as $\mathcal{M}_{\mathbf{x}_{m}\rightarrow f_{\mathbf{y}_{m}}}(\mathbf{x}_{m})=\mathcal{CN}(\mathbf{x}_{m};\mathbf{x}_{m,\mathsf{A}}^{\mathtt{pri}},v_{m,\mathsf{A}}^{\mathtt{pri}}\mathbf{I}_{NK})$. The belief of $\mathbf{x}_{m}$ on the factor node $p(\mathbf{y}_{m}|\mathbf{x}_{m})$ is expressed as
@@
The message from the factor node $p(\mathbf{y}_{m}|\mathbf{x}_{m})$ to the variable node $\mathbf{x}_{m}$, also known as the extrinsic message, is calculated based on the EP algorithm [49] as
@@ III-B Module B: Channel and Device Activity Denoisers
As shown in Fig. 2, the denoiser of $\mathbf{X}_{k}=\alpha_{k}\mathbf{H}_{k}$ can be further divided into two sub-modules: the denoiser of $\mathbf{H}_{k}$ and the denoiser of $\alpha_{k}$. The outputs from these two sub-modules are then combined at the factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ to produce the denoised $\mathbf{X}_{k}$. In this subsection, we formally establish the equivalence between belief calculation at the factor node $p(\mathbf{H}_{k})$ and MMSE denoising of $\mathbf{H}_{k}$ under AWGN observations. This equivalence plays a crucial role in the integration of score-based generative priors for efficient belief calculation.
Firstly, we derive the message from the variable node $\mathbf{X}_{k}$ to the factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$, denoted as $\mathcal{M}_{\mathbf{X}_{k}\rightarrow f_{\mathbf{X}_{k}}}(\mathbf{X}_{k})$, by examining the relationships in vector-matrix reorganization. We marginalize the message $\mathcal{M}_{f_{\mathbf{y}_{m}}\rightarrow\mathbf{x}_{m}}(\mathbf{x}_{m})=\mathcal{CN}\big(\mathbf{x}_{m};\mathbf{x}_{m,\mathsf{B}}^{\mathtt{pri}},v_{m,\mathsf{B}}^{\mathtt{pri}}\mathbf{I}_{NK}\big)$ from module A to obtain
@@
where $\mathbf{x}_{km}\triangleq\left[x_{k1m},\dots,x_{kNm}\right]^{\sf T}\in\mathbb{C}^{N}$ and $\mathbf{x}_{km,\mathsf{B}}^{\mathtt{pri}}\triangleq\big[x_{k1m,\mathsf{B}}^{\mathtt{pri}},\dots,x_{kNm,\mathsf{B}}^{\mathtt{pri}}\big]^{\sf T}\in\mathbb{C}^{N}$ represent the $k$-th blocks in $\mathbf{x}_{m}$ and $\mathbf{x}_{m,\mathsf{B}}^{\mathtt{pri}}$, respectively, i.e., $\mathbf{x}_{m}=\left[\mathbf{x}_{1m}^{\sf H},\dots,\mathbf{x}_{Km}^{\sf H}\right]^{\sf H}$ and $\mathbf{x}_{m,\mathsf{B}}^{\mathtt{pri}}=\big[(\mathbf{x}_{1m,\mathsf{B}}^{\mathtt{pri}})^{\sf H},\dots,(\mathbf{x}_{Km,\mathsf{B}}^{\mathtt{pri}})^{\sf H}\big]^{\sf H}$. After vector-matrix reorganization, the message from the variable node $\mathbf{X}_{k}$ to the factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ manifests as
@@ III-B1 Denoiser of $\mathbf{H}_{k}$
The message from the factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ to the variable node $\mathbf{H}_{k}$ is computed following the sum-product rule [48]:
@@
where $\mathbf{h}_{km}\triangleq\left[h_{k1m},\dots,h_{kNm}\right]^{\sf T}\in\mathbb{C}^{N}$ represents the $m$-th column of $\mathbf{H}_{k}$, i.e., $\mathbf{H}_{k}=\left[\mathbf{h}_{k1},\dots,\mathbf{h}_{kM}\right]$. The first term in (15) corresponds to the case that the $k$-th device is inactive and is a constant with respect to $\mathbf{H}_{k}$. As a result, $\mathcal{M}_{f_{\mathbf{X}_{k}}\rightarrow\mathbf{H}_{k}}(\mathbf{H}_{k})$ cannot be properly normalized to a valid pdf when this term is retained. To obtain a tractable message for channel denoising, we approximate $\mathcal{M}_{f_{\mathbf{X}_{k}}\rightarrow\mathbf{H}_{k}}(\mathbf{H}_{k})$ by conditioning on the active case and forwarding only the second term to the channel denoiser. The activity uncertainty will be handled separately through the Bernoulli update in (30). The approximated Gaussian message reads
@@
Based on the update rules of EP [49], the belief of $\mathbf{H}_{k}$ on the factor node $p(\mathbf{H}_{k})$ is expressed as
@@
$~~~~~~~~~~~~~~~~~~~~~~\times\prod_{m}\mathcal{CN}\big(\mathbf{h}_{km};\mathbf{h}_{km}^{\mathtt{pri}},\tau_{m}^{\mathtt{pri}}\mathbf{I}_{N}\big),$ (19)
@@
State evolution (SE) analysis [51] has shown that the error vector $\mathbf{h}_{km}^{\mathtt{pri}}-\mathbf{h}_{km}$ is i.i.d. Gaussian in the large system limit. In other words, module A decouples the linear mixing model (3) into a series of independent AWGN observations of the ground truth, i.e.,
@@
where $\mathbf{H}_{k}^{\mathtt{post}}\triangleq\left[\mathbf{h}_{k1}^{\mathtt{post}},\dots,\mathbf{h}_{kM}^{\mathtt{post}}\right]$. To summarize, the belief calculation amounts to combining the Gaussian pseudo-likelihood $p(\mathbf{H}_{k}^{\mathtt{pri}}|\mathbf{H}_{k})$ with the true prior $p(\mathbf{H}_{k}).$ This operation is mathematically equivalent to MMSE denoising under AWGN observations: the pseudo-observation from Module A provides a noisy estimate of the channel, and the prior serves as the denoiser to refine this estimate. The posterior mean and variance in (22) and (23) are, however, by no means easy to calculate due to the complicated and high-dimensional integrals. We will detail the method for obtaining the posterior message in Sections IV and V.
The extrinsic message from the variable node $\mathbf{H}_{k}$ to the factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ is calculated as
@@ III-B2 Denoiser of $\alpha_{k}$
The belief of $\alpha_{k}$ on the factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ is expressed as
@@
is the posterior probability of activity for the $k$-th device. We set a threshold $\lambda^{\mathtt{thr}}$ to facilitate the binary decision of device activity, i.e.,
@@ III-B3 Denoiser of $\mathbf{X}_{k}$
Based on the update rules of EP [49], the belief of $\mathbf{X}_{k}$ on the factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ is expressed as
@@
$~~~~~~~~~~~~~~~~~~~\times\mathcal{M}_{\mathbf{H}_{k}\rightarrow f_{\mathbf{X}_{k}}}(\mathbf{H}_{k})p(\alpha_{k})\bigg]$
@@
While per-element or full covariance estimation could, in principle, capture heterogeneous confidence across users and subcarriers, such estimates are highly unreliable with limited pilot observations. Scalar averaging, as adopted in (38), is a standard design choice to stabilize message-passing updates [26, 50].
Finally, we calculate the message from the factor node $p(\mathbf{X}_{k}|\mathbf{H}_{k},\alpha_{k})$ to the variable node $\mathbf{X}_{k}$ as
@@ IV-A Bridging Score Function and MMSE Denoising
where $\nabla_{\mathbf{H}_{k}^{\mathtt{pri}}}\log p(\mathbf{H}_{k}^{\mathtt{pri}})\in\mathbb{C}^{N\times M}$ is the gradient of the log-density function, commonly known as the (first-order) score function. Tweedie’s formula bridges MMSE denoising and the score of the perturbed data distribution through a simple equation, eliminating the need for calculating high-dimensional integrals. Specifically, the posterior mean can be viewed as the sum of the maximum likelihood estimator and a Bayes correction term. The formula can also be framed as “denoised result = noisy data - predicted noise” from a denoising perspective, in which the predicted noise corresponds to $\tau^{\mathtt{pri}}\nabla_{\mathbf{H}_{k}^{\mathtt{pri}}}\log p(\mathbf{H}_{k}^{\mathtt{pri}})$.
@@
where $\nabla_{\mathbf{H}_{k}^{\mathtt{pri}}}^{2}\log p(\mathbf{H}_{k}^{\mathtt{pri}})\in\mathbb{C}^{NM\times NM}$ is the Hessian of the log-density function, referred to as the second-order score function. The first- and second-order Tweedie’s formulas are well-known results that can be proven using simple integration tricks [52, Appendix B] or by applying the properties of exponential families [53, Appendix A].
@@ IV-B First-Order Score Matching
[…]
@@
[…]
@@
A common choice for the weighting factor is $\lambda_{1}(\tau_{i})=\tau_{i}$ [15], which empirically balances the loss term $\ell_{1}(\tau_{i};\boldsymbol{\theta})$ for different values of $\tau_{i}$.
@@ IV-C Second-Order Score Matching
The second-order score matching objective can also be expressed in an equivalent denoising score matching formulation as
@@
[…]
@@
where $\lambda_{2}(\tau_{i})$ is the weighting factor depending on $\tau_{i}$. Following [31], we choose $\lambda_{2}(\tau_{i})=\tau_{i}^{2}$.
@@
Figure 3: Diagram of the STMP-JADCE algorithm. The real and imaginary parts of the wireless channel matrix are concatenated as two feature channels (analogous to the RGB channels in images) when fed into the score networks. For the $k$-the device, the input tensor is constructed by stacking the real and imaginary parts of $\mathbf{H}_{k}$ into a tensor of size $(N,M,2)$. Across multiple devices, these tensors are further batched along the user dimension. In this figure, the first and the second devices are active, whereas the $K$-th device is inactive. Therefore, the input to the score networks takes the form of $\mathbf{H}_{k}^{\mathtt{pri}}=\mathbf{H}_{k}+\mathbf{W}_{k}$ for $k=1,2$, while $\mathbf{H}_{K}^{\mathtt{pri}}$ is a random noise. Note that, for brevity, we omitted the skip connections (dependences) from the prior messages to both the the posterior and extrinsic messages. Please refer to eqns. (55), (56), (26), and (27) for the rigorous relationships.
@@ V-A Overall Algorithm
where $s_{2,nm}(\mathbf{H}_{k}^{\mathtt{pri}},\tau^{\mathtt{pri}};\boldsymbol{\phi})$ denotes the ((n1)M+m)((n-1)M+m)-th diagonal elements of $\mathbf{S}_{2}\big(\mathbf{H}_{k}^{\mathtt{pri}},\tau^{\mathtt{pri}};\boldsymbol{\phi}\big)$.
Fig. 3 demonstrates the integration of the two score networks into the TMP framework, resulting in the STMP-JADCE algorithm. STMP-JADCE features a hybrid structure in which the channel denoiser is realized by score networks, while the device activity denoiser and the linear estimator are derived analytically. In this way, our method can be easily applied across different device activity levels, pilot patterns, and SNRs without the need to retain the score networks. Moreover, the potential mismatch between the chosen prior of the device activity and the ground truth can be rectified by incorporating the expectation-maximization (EM) update [54]. Essentially, the development of STMP-JADCE highlights a broader philosophical shift: whereas conventional Bayesian inference in wireless communications has long relied on manually specified and typically oversimplified priors for tractability, our framework embraces a data-driven paradigm in which the channel prior is learned directly from data. This data-driven Bayesian inference not only resolves the mismatch between hand-crafted priors and real channel statistics, but also opens the door to more principled integration of modern generative models into inference algorithms. We discuss the engineering tricks used in STMP-JADCE in the following.